OpenStudy (anonymous):

Which value for x makes the sentence true? 0.02x = 0.0994 0.1194 0.001988 0.0794 4.97

OpenStudy (alexandervonhumboldt2):

divide both sides by 0.02

OpenStudy (anonymous):

need help fast

OpenStudy (alexandervonhumboldt2):

0.02x/0.02=0.0994/0.02 x=4.97

OpenStudy (anonymous):

Which value for x makes the sentence true? 2/3+i=9/10

OpenStudy (alexandervonhumboldt2):

x or i?

OpenStudy (anonymous):

i

OpenStudy (alexandervonhumboldt2):

ok so subtarct 2/3 from both sides: 2/3-2/3+i=9/10-2/3 i=9/10-2/3 equal the denominators: i=27/30-20/30 i=7/30 thus i=7/30

OpenStudy (anonymous):

need help fast

OpenStudy (anonymous):

Which value for x makes the sentence true? 4,361 + x = 8,477